OpenStudy (anonymous):

Solve the exponential equation. Express the solution in terms of logarithms. 7^(4x-2) = 6^(x-2)

OpenStudy (campbell_st):

take the log of both sides... doesn't matter if its base e or base 10 \[\log (7^{4x - 2} = \log(6^{x - 2})\] then by using log laws \[(4x -2)\log(7) = (x -1) \log(6)\] expand anc collect like terms 4xlog(7) - x log(6) = 2log(7) - log(6) x(4log(7) - log(6)) = 2log(7) - log(6) \[x = \frac{2\log(7) - \log(6)}{2\log(7) - \log(6)}\] which can be simplified further to \[x = \frac{\log(\frac{7^4}{6})}{\log(\frac{7^2}{6})}\]

OpenStudy (anonymous):

perfect thank you. awesome step by step process

OpenStudy (anonymous):

I believe that this problem was solved in reverse. The answer should be the same as given only with the numerator and denominator swapped. so Log(7^4/6) should be in the denominator.

OpenStudy (anonymous):

\[x = \log(7^2/6)/\log(7^4/6)\]